# Cellular, Cell-less, and Everything in Between: A Unified Framework for Utility Region Analysis in Wireless Networks

### v2

@@ II-A Notation and standard definitions
Let $\|\cdot\|$ be a monotone norm on ${\mathbb{R}}^{N}$, and define $\mathcal{N}:=\{1,\ldots,N\}$. We say that a (standard interference) mapping $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}:{\bm{x}}\mapsto(t_{1}({\bm{x}}),\ldots,t_{N}({\bm{x}}))$ satisfies hypothesis ($T{T_{\|\cdot\|}}$) if there exist sets $(\mathcal{Y}_{n})_{n\in\mathcal{N}}$, positive scalars $(u_{y,n})_{n\in\mathcal{N},y\in\mathcal{Y}_{n}}$ bounded away from zero, and general interference functions $(h_{y,n})_{n\in\mathcal{N},y\in\mathcal{Y}_{n}}$, such that the coordinate standard interference functions can be expressed as
$(\forall n\in\mathcal{N})(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})~t_{n}({\bm{x}}):=\inf_{y\in\mathcal{Y}_{n}}(h_{y,n}({\bm{x}})+{u}_{y,n}).$
In this case, we denote by $T_{\|\cdot\|}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ the general interference mapping given by
$T_{\|\cdot\|}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{++}^{N}:{\bm{x}}\mapsto({(t_{1})_{\|\cdot\|}}({\bm{x}}),\ldots,{(t_{N})_{\|\cdot\|}}({\bm{x}})),$ (8)
where, for each $n\in\mathcal{N}$, the general interference function ${(t_{n})_{\|\cdot\|}}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ is as defined in (7); i.e.,
$(\forall n\in\mathcal{N})(\forall{\bm{x}}\in{\mathbb{R}}^{N}_{+})~(t_{n})_{\|\cdot\|}({\bm{x}}):=\inf_{y\in\mathcal{Y}_{n}}(h_{y,n}({\bm{x}})+{u}_{y,n}~\|{\bm{x}}\|).$
Remark 1.
The mappings $T$ and $T_{\|\cdot\|}$} in Definition 4 are respectively, a standard interference mapping and a general interference mapping, so their restrictions to the positive cone ${\mathbb{R}}_{++}^{N}$ are necessarily continuous, and these mappings restricted to the domain ${\mathbb{R}}_{++}^{N}$ can be continuously extended to the entire nonnegative cone ${\mathbb{R}}_{+}^{N}$ [33, Theorem 5.1.5]. Therefore, to avoid unnecessary technical digressions, we tacitly assume that both $T$ and $T_{\|\cdot\|}$} are continuous on ${\mathbb{R}}_{+}^{N}$.
We now proceed to show that important properties of conditional eigenvalues – which are typically related to SINR values or rates in wireless network models [1, 23] – associated with a monotone norm $\|\cdot\|$ and a mapping $T$ satisfying hypothesis ($T{T_{\|\cdot\|}}$) can be deduced from the spectral radius $\rho(T_{\|\cdot\|})$ and the corresponding eigenvector of $T_{\|\cdot\|}$ in Definition 4. The mapping $T_{\|\cdot\|}$ unifies $T$ and the norm $\|\cdot\|$ into a single concept that is often more convenient for analysis. The next proposition also shows that $T_{\|\cdot\|}$ possesses a unique eigenvector (up to a scaling factor) associated with the spectral radius, which is a special result considering that we deal with possibly nonlinear mappings.
@@
Given a monotone norm $\|\cdot\|$ on ${\mathbb{R}}^{N}$, assume that $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ satisfies hypothesis ($T{T_{\|\cdot\|}}$). Let ${\bm{x}}^{\star}\in{\mathbb{R}}_{++}^{N}$ and $\lambda^{\star}>0$ be, respectively, the conditional eigenvector and the conditional eigenvalue associated with $T$ and $\|\cdot\|$. Then each of the following holds:
@@
the set of eigenvectors of $T_{\|\cdot\|}$} associated with the spectral radius $\rho({T_{\|\cdot\|}})$ is given by $\{\alpha{\bm{x}}^{\star}\in{\mathbb{R}}_{++}^{N}\mid\alpha>0\}$.
@@
where (a) follows from the fact that $T_{\|\cdot\|}$ is a general interference mapping (Lemma 1), and, hence, positively homogeneous; (b) follows from $\|{\bm{w}}\|=1$ and the definitions of $T$ and $T_{\|\cdot\|}$}; and (c) follows from positivity of $T$. The above shows that $\bm{w}$} and $\rho({T_{\|\cdot\|}})$ are, respectively, a conditional eigenvector and a conditional eigenvalue associated with $T$ and the monotone norm $\|\cdot\|$. Uniqueness of the conditional eigenvalue and eigenvector [1] (Fact 6(i) in the Supplemental Material) implies that $\rho({T_{\|\cdot\|}})=\lambda^{\star}$ and ${\bm{w}}=(1/\|{\bm{z}}\|){\bm{z}}={\bm{x}}^{\star}$, which completes the proof of parts (i) and (ii).
@@
In the remainder of this section, for a given monotone norm $\|\cdot\|$ and a mapping $T$ satisfying hypothesis ($T{T_{\|\cdot\|}}$), we establish conditions under which the function ${\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto\rho({\mathrm{diag}(\bm{x})}T_{\|\cdot\|})$i.e., the function mapping a given ${\bm{x}}\in{\mathbb{R}}_{+}^{N}$ to the spectral radius of the general interference mapping ${\mathbb{R}}^{N}_{+}\to{\mathbb{R}}^{N}_{+}:{\bm{u}}\mapsto\mathrm{diag}({\bm{x}})({T_{\|\cdot\|}}({\bm{u}}))$ – satisfies all properties of a norm in the nonnegative cone ${\mathbb{R}}^{N}_{+}$. We then extend this cone-defined norm to the entire space $\mathbb{R}^{N}$. In the next section, we show that this extension enables us to analyze interference models in wireless networks using concise notation that immediately reveals high-level properties of the feasible SINR and rate regions. For instance, by simply glancing at the definition of the feasible SINR region in (28) in Sect. III-B below, readers can infer from the notation that this region resembles a ball defined by some norm, making it both convex and compact, among other intuitive properties. Remarkably, this insight is achieved without requiring any knowledge of the specific details of the interference model, the wireless technology, or the norm under consideration.
@@
Let $\|\cdot\|$ be a monotone norm and $T:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}^{N}_{++}$ satisfy hypothesis ($T{T_{\|\cdot\|}}$). Consider the function $g_{T_{\|\cdot\|}}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto\rho(\mathrm{diag}({\bm{x}})T_{\|\cdot\|})$. Then each of the following holds:
@@
(iv) For a matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}$, denote by $\|{\bm{M}}\|$ the matrix operator norm induced by the vector norm $\|\cdot\|$. Let $\mathcal{P}$ be the compact set given by $\mathcal{P}:=\{{\bm{x}}\in{\mathbb{R}}_{+}^{N}\mid\|{\bm{x}}\|=1\}$. Since $T_{\|\cdot\|}$} is a general interference mapping, and hence continuous (see Remark 1), the extreme-value theorem shows that there exists $B\in{\mathbb{R}}_{+}$ such that $B=\sup_{{\bm{p}}\in\mathcal{P}}\|{T_{\|\cdot\|}}({\bm{p}})\|$, and the supremum is attained at some ${\bm{p}}^{\star}\in\mathcal{P}$. As a result, if a sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ in ${\mathbb{R}}_{+}^{N}$ converges to ${\bm{x}}^{\star}\in{\mathbb{R}}_{+}^{N}$, we can use the standard property $(\forall{\bm{x}}\in{\mathbb{R}}^{N})(\forall{\bm{M}}\in{\mathbb{R}}^{N\times N})\|{\bm{Mx}}\|\leq\|{\bm{M}}\|~\|{\bm{x}}\|$ of matrix operator norms to deduce
@@
We can now apply [29, Theorem 3.1(2)] (Fact 8 in the Supplemental Material) to deduce:
@@
The function $g_{T_{\|\cdot\|}}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ defined in Proposition 2 satisfies almost all the properties of a norm in the nonnegative cone ${\mathbb{R}}_{+}^{N}$. The only exception is the triangle inequality. This missing property is automatically satisfied in the nonnegative cone if $g_{T_{\|\cdot\|}}$ is known to be convex only in the positive cone ${\mathbb{R}}_{++}^{N}$, in which case we can extend $g_{T_{\|\cdot\|}}$ to the entire domain ${\mathbb{R}}^{N}$ while keeping the properties of a norm:
@@
Let $\|\cdot\|$ be a monotone norm on ${\mathbb{R}}^{N}$, $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ satisfy hypothesis ($T{T_{\|\cdot\|}}$), and $g_{T_{\|\cdot\|}}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ be as defined in Proposition 2. Assume that $g_{T_{\|\cdot\|}}$ is a convex function in ${\mathbb{R}}_{++}^{N}$. Then $g_{T_{\|\cdot\|}}$ is also convex in ${\mathbb{R}}_{+}^{N}$. Furthermore, for $C_{T_{\|\cdot\|}}:=\{{\bm{x}}\in{\mathbb{R}}^{N}_{+}~\mid g_{T_{\|\cdot\|}}({\bm{x}})\leq 1\}$ and $S_{T_{\|\cdot\|}}=\mathrm{conv}(C_{T_{\|\cdot\|}}\cup-C_{T_{\|\cdot\|}})$, where $C_{T_{\|\cdot\|}}:=\{{\bm{x}}\in{\mathbb{R}}^{N}\mid-{\bm{x}}\in C_{T_{\|\cdot\|}}\}$ and $\mathrm{conv}(\cdot)$ is the convex hull of a set, the Minkowski or gauge functional of $S_{T_{\|\cdot\|}}$, defined by
@@
Here we use extensively the properties of $g_{T_{\|\cdot\|}}$ in Proposition 2(i)-(iv). The proof that convexity on the interior of the domain of a everywhere continuous function implies convexity on the whole domain is standard, but we show the main steps for completeness. Assume that $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ and $({\bm{y}}_{n})_{n\in{\mathbb{N}}}$ are two sequences in ${\mathbb{R}}_{++}^{N}$ converging in the normed vector space $({\mathbb{R}}^{N},\|\cdot\|)$ to, respectively, the vectors ${\bm{x}}\in{\mathbb{R}}_{+}^{N}$ and ${\bm{y}}\in{\mathbb{R}}_{+}^{N}$. Given $\alpha\in~]0,1[$, convexity of $g_{T_{\|\cdot\|}}$ on ${\mathbb{R}}_{++}^{N}$ implies that $(\forall n\in{\mathbb{N}})~g_{T_{\|\cdot\|}}(\alpha{\bm{x}}_{n}+(1-\alpha){\bm{y}}_{n})\leq\alpha g_{T_{\|\cdot\|}}({\bm{x}}_{n})+(1-\alpha)g_{T_{\|\cdot\|}}({\bm{y}}_{n})$. Passing to the limit as $n\to\infty$ and recalling that $g_{T_{\|\cdot\|}}$ is continuous on ${\mathbb{R}}^{N}_{+}$, we deduce $g_{T_{\|\cdot\|}}(\alpha{\bm{x}}+(1-\alpha){\bm{y}})\leq\alpha g_{T_{\|\cdot\|}}({\bm{x}})+(1-\alpha)g_{T_{\|\cdot\|}}({\bm{y}})$, which establishes convexity of $g_{T_{\|\cdot\|}}$ on ${\mathbb{R}}_{+}^{N}$. In particular, setting $\alpha$ to $\alpha=1/2$ and using homogeneity of $g_{T_{\|\cdot\|}}$, we deduce $\alpha g_{T_{\|\cdot\|}}({\bm{x}}+{\bm{y}})\leq\alpha g_{T_{\|\cdot\|}}({\bm{x}})+\alpha g_{T_{\|\cdot\|}}({\bm{y}})$, which proves that $g_{T_{\|\cdot\|}}$ satisfies the triangle inequality on ${\mathbb{R}}_{+}^{N}$. Therefore, $g_{T_{\|\cdot\|}}$ satisfies all properties of a norm when restricted to vectors in the closed cone ${\mathbb{R}}_{+}^{N}$, so the set $C_{T_{\|\cdot\|}}:=\{{\bm{x}}\in{\mathbb{R}}^{N}_{+}~\mid g_{T_{\|\cdot\|}}({\bm{x}})\leq 1\}$ is a compact convex set with nonempty interior. Furthermore, $C_{T_{\|\cdot\|}}$ is downward comprehensive on ${\mathbb{R}}_{+}^{N}$ as a consequence of Proposition 2(iii). All the properties of $C_{T_{\|\cdot\|}}$ we have proved enable us to apply [35, Proposition 2] to conclude that the Minkowski functional in (10) is indeed a monotone norm, and it satisfies
@@
The result in Proposition 3 provides a foundation for a natural definition of a norm induced by a general interference mapping:
@@
We say that a general interference mapping $G:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ is norm inducing if the following properties hold:
  • (i)
    there exist a monotone norm $\|\cdot\|$ and a standard interference mapping $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ satisfying hypothesis ($T{T_{\|\cdot\|}}$) such that $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})~G({\bm{x}})={T_{\|\cdot\|}}({\bm{x}})$; and
  • (ii)
    the function $g_{G}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ defined by $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})~g_{G}({\bm{x}}):=\rho(\mathrm{diag}({\bm{x}})G)=\rho(\mathrm{diag}({\bm{x}}){T_{\|\cdot\|}})$ is convex on ${\mathbb{R}}_{++}^{N}$.
In this case, based on Proposition 3, we call the norm $\|\cdot\|_{G}:=\|\cdot\|_{T_{\|\cdot\|}}$ in (10) the norm induced by $G$.
Our primary focus is on constructing norm-inducing general interference mappings derived from an explicitly given standard interference mapping and a monotone norm. However, in certain instances, it is straightforward to extract a suitable standard interference mapping and a monotone norm from a general interference mapping $G$ to verify that $G$ is norm inducing, as illustrated in the next example, which is common in wireless networks – see, for example, the set in (23) in Sect. III-A. The proof of the statement in the next example is shown in the Supplementary Material.

### v3

@@ II-A Notation and standard definitions
Let $\mathcal{N}:=\{1,\ldots,N\}$. We say that a standard interference mapping $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}:{\bm{x}}\mapsto(t_{1}({\bm{x}}),\ldots,t_{N}({\bm{x}}))$ and a general interference mapping $T_{\|\cdot\|}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{+}^{N}:{\bm{x}}\mapsto({(t_{1})_{\|\cdot\|}}({\bm{x}}),\ldots,{(t_{N})_{\|\cdot\|}}({\bm{x}}))$ satisfy hypothesis ($T{T_{\|\cdot\|}}$) for a monotone norm $\|\cdot\|$ on ${\mathbb{R}}^{N}$ if there exist sets $(\mathcal{Y}_{n})_{n\in\mathcal{N}}$, a family $(u_{y,n})_{n\in\mathcal{N},y\in\mathcal{Y}_{n}}$ of positive scalars bounded away from zero, and general interference functions $(h_{y,n})_{n\in\mathcal{N},y\in\mathcal{Y}_{n}}$ such that the coordinate functions $t_{n}$ and $(t_{n})_{\|\cdot\|}$ for each $n\in\mathcal{N}$ can be expressed as
$t_{n}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}:{\bm{x}}\mapsto\inf_{y\in\mathcal{Y}_{n}}(h_{y,n}({\bm{x}})+{u}_{y,n})$ (11)
and
$(t_{n})_{\|\cdot\|}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto\inf_{y\in\mathcal{Y}_{n}}(h_{y,n}({\bm{x}})+{u}_{y,n}~\|{\bm{x}}\|).$ (12)
As a simple illustration of mappings satisfying hypothesis ($T{T_{\|\cdot\|}}$) for a given monotone norm $\|\cdot\|$, we have:
Example 2.
Under the stated assumptions on $\bm{M}$}, ${\bm{u}}=:[u_{1},\ldots,u_{N}]^{t}$, $\mathcal{N}:=\{1,\ldots,N\}$, and the monotone norm $\|\cdot\|$ in Example 1, we verify that $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}:{\bm{x}}\mapsto{\bm{Mx}}+{\bm{u}}$ and $T_{\|\cdot\|}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}:{\bm{x}}\mapsto{\bm{Mx}}+{\bm{u}}\|{\bm{x}}\|$ satisfy hypothesis $(T,{T_{\|\cdot\|}})$. More specifically, the $n$th coordinate functions $(t_{n})_{n\in\mathcal{N}}$ and $(t_{n})_{\|\cdot\|}$ of, respectively, $T$ and $T_{\|\cdot\|}$ are given by $t_{n}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{++}:{\bm{x}}\mapsto h_{n}({\bm{x}})+{u}_{n}$ and $(t_{n})_{\|\cdot\|}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto h_{n}({\bm{x}})+{u}_{n}\|{\bm{x}}\|$, where $(\forall n\in\mathcal{N})~h_{n}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto{\bm{m}}_{n}^{t}{\bm{x}}$ is a general interference function and ${\bm{m}}_{n}^{t}$ is a row vector corresponding to the $n$th row of the matrix $\bm{M}$}.
Let $T$ and $T_{\|\cdot\|}$} satisfy hypothesis $(T,{T_{\|\cdot\|}})$ for a given norm $\|\cdot\|$. We now proceed to show that important properties of conditional eigenvalues – which are typically related to SINR values or rates in wireless network models [29, 9] – associated with the monotone norm and the mapping $T$ can be deduced from the spectral radius $\rho(T_{\|\cdot\|})$ and the corresponding eigenvector of $T_{\|\cdot\|}$. The mapping $T_{\|\cdot\|}$ unifies $T$ and the norm $\|\cdot\|$ into a single concept that is often more convenient for analysis. The next proposition also shows that $T_{\|\cdot\|}$ possesses a unique eigenvector (up to a scaling factor) associated with the spectral radius, which is a special result considering that we deal with possibly nonlinear mappings.
@@
Given a monotone norm $\|\cdot\|$ on ${\mathbb{R}}^{N}$, assume that $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ and ${T_{\|\cdot\|}}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ satisfy hypothesis ($T{T_{\|\cdot\|}}$). Let ${\bm{x}}^{\star}\in{\mathbb{R}}_{++}^{N}$ and $\lambda^{\star}>0$ be, respectively, the conditional eigenvector and the conditional eigenvalue associated with $T$ and $\|\cdot\|$ (see Fact 6(i) in the Supplemental Material). Then each of the following holds:
@@
the set of eigenvectors of $T_{\|\cdot\|}$} associated with the spectral radius $\rho({T_{\|\cdot\|}})$ is the cone given by $\{\alpha{\bm{x}}^{\star}\in{\mathbb{R}}_{++}^{N}\mid\alpha>0\}$.
@@
where (a) follows from the fact that $T_{\|\cdot\|}$ is a general interference mapping (Lemma 1), and, hence, positively homogeneous; (b) follows from $\|{\bm{w}}\|=1$ and the definitions of $T$ and $T_{\|\cdot\|}$}; and (c) follows from positivity of $T$. The above shows that $\bm{w}$} and $\rho({T_{\|\cdot\|}})$ are, respectively, a conditional eigenvector and a conditional eigenvalue associated with $T$ and the monotone norm $\|\cdot\|$. Uniqueness of the conditional eigenvalue and eigenvector (Fact 6(i) in the Supplemental Material) implies that $\rho({T_{\|\cdot\|}})=\lambda^{\star}$ and ${\bm{w}}=(1/\|{\bm{z}}\|){\bm{z}}={\bm{x}}^{\star}$, which completes the proof of parts (i) and (ii).
@@
We are now in a position to make precise the claims stated in Remark 1.
Example 3.
Consider the mappings $T$ and $T_{\|\cdot\|}$} defined in Example 2 under the assumptions stated on $\bm{M}$}, $\bm{u}$}, $x_{\mathrm{max}}$, and $\|\cdot\|$ in Example 1. We verify from [26, Eq. (10)] that a solution ${\bm{x}}^{\star}\in{\mathbb{R}}_{++}^{N}$ to Problem (8) satisfies $T({\bm{x}}^{\star})=(1/t^{\star}){\bm{x}}^{\star}$ and $\|{\bm{x}}^{\star}\|=1$, where $t^{\star}>0$ is the optimal objective of Problem (8). As a result, $\bm{x}$$^{\star}$ and $1/t^{\star}$ are, respectively, the conditional eigenvector and eigenvalue corresponding to the mapping $T$ and the monotone norm $\|\cdot\|$ in Example 1. The claims in Remark 1 now follow immediately from Proposition 1.
In the remainder of this section, for a given monotone norm $\|\cdot\|$ and mappings $T$ and $T_{\|\cdot\|}$} satisfying hypothesis ($T{T_{\|\cdot\|}}$), we establish conditions under which the function $g_{{T_{\|\cdot\|}}}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto\rho({\mathrm{diag}(\bm{x})}T_{\|\cdot\|})$i.e., the function mapping a given ${\bm{x}}\in{\mathbb{R}}_{+}^{N}$ to the spectral radius of the general interference mapping ${\mathbb{R}}^{N}_{+}\to{\mathbb{R}}^{N}_{+}:{\bm{y}}\mapsto\mathrm{diag}({\bm{x}})({T_{\|\cdot\|}}({\bm{y}}))=(\mathrm{diag}({\bm{x}})\circ{T_{\|\cdot\|}})({\bm{y}})$ [^3] Here, given ${\bm{x}}\in{\mathbb{R}}_{+}^{N}$, the matrix $\mathrm{diag}({\bm{x}})\in{\mathbb{R}}_{+}^{N\times N}$ should be interpreted as the linear operator $\mathrm{diag}({\bm{x}}):{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}:{\bm{y}}\mapsto\mathrm{diag}({\bm{x}}){\bm{y}}$. satisfies all properties of a norm on the nonnegative cone ${\mathbb{R}}^{N}_{+}$. We then extend this cone-defined norm to the entire space $\mathbb{R}^{N}$. In the next section, we show that this extension enables us to analyze interference models in wireless networks using concise notation that immediately reveals high-level properties of the feasible SINR and rate regions. For instance, by simply glancing at the definition of the feasible SINR region in (33) in Sect. III-C below, readers can infer from the notation that this region resembles a ball defined by some norm, making it both convex and compact, among other intuitive properties. Remarkably, this insight is achieved without requiring any knowledge of the specific details of the interference model, the wireless technology, or the norm under consideration.
@@
Let $\|\cdot\|$ be a monotone norm and $T:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}^{N}_{++}$ and ${T_{\|\cdot\|}}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}^{N}_{+}$ satisfy hypothesis ($T{T_{\|\cdot\|}}$). Consider the function $g_{T_{\|\cdot\|}}:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto\rho(\mathrm{diag}({\bm{x}})T_{\|\cdot\|})$. Then each of the following holds:
@@
(iv) For a matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}$, denote by $\|{\bm{M}}\|$ the matrix operator norm induced by the vector norm $\|\cdot\|$. Let $\mathcal{P}$ be the compact set given by $\mathcal{P}:=\{{\bm{x}}\in{\mathbb{R}}_{+}^{N}\mid\|{\bm{x}}\|=1\}$. Since $T_{\|\cdot\|}$} is a general interference mapping, and hence continuous according to our definition (see the discussion below Definition 1), the extreme-value theorem shows that there exists $B\in{\mathbb{R}}_{+}$ such that $B=\sup_{{\bm{p}}\in\mathcal{P}}\|{T_{\|\cdot\|}}({\bm{p}})\|$, and the supremum is attained at some ${\bm{p}}^{\star}\in\mathcal{P}$. As a result, if a sequence $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ in ${\mathbb{R}}_{+}^{N}$ converges to ${\bm{x}}^{\star}\in{\mathbb{R}}_{+}^{N}$, we can use the standard property $(\forall{\bm{x}}\in{\mathbb{R}}^{N})(\forall{\bm{M}}\in{\mathbb{R}}^{N\times N})\|{\bm{Mx}}\|\leq\|{\bm{M}}\|~\|{\bm{x}}\|$ of matrix operator norms to deduce
@@
We can now apply Fact 8 in the Supplemental Material to deduce:
@@
The function $g_{T_{\|\cdot\|}}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ defined in Proposition 2 satisfies almost all the properties of a norm on the nonnegative cone ${\mathbb{R}}_{+}^{N}$. The only exception is the triangle inequality. This missing property is automatically satisfied on the nonnegative cone if $g_{T_{\|\cdot\|}}$ is known to be convex only in the positive cone ${\mathbb{R}}_{++}^{N}$, in which case we can extend $g_{T_{\|\cdot\|}}$ to the entire domain ${\mathbb{R}}^{N}$ while keeping the properties of a norm:
@@
Let $\|\cdot\|$ be a monotone norm on ${\mathbb{R}}^{N}$, $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ and ${T_{\|\cdot\|}}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ satisfy hypothesis ($T{T_{\|\cdot\|}}$), and $g_{T_{\|\cdot\|}}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ be as defined in Proposition 2. Assume that $g_{T_{\|\cdot\|}}$ is a convex function in ${\mathbb{R}}_{++}^{N}$. Then $g_{T_{\|\cdot\|}}$ is also convex in ${\mathbb{R}}_{+}^{N}$. Furthermore, for $C_{T_{\|\cdot\|}}:=\{{\bm{x}}\in{\mathbb{R}}^{N}_{+}~\mid g_{T_{\|\cdot\|}}({\bm{x}})\leq 1\}$ and $S_{T_{\|\cdot\|}}=\mathrm{conv}(C_{T_{\|\cdot\|}}\cup-C_{T_{\|\cdot\|}})$, where $C_{T_{\|\cdot\|}}:=\{{\bm{x}}\in{\mathbb{R}}^{N}\mid-{\bm{x}}\in C_{T_{\|\cdot\|}}\}$ and $\mathrm{conv}(\cdot)$ is the convex hull of a set, the Minkowski or gauge functional of $S_{T_{\|\cdot\|}}$, defined by
@@
Here we use extensively the properties of $g_{T_{\|\cdot\|}}$ in Proposition 2(i)-(iv). The proof that convexity on the interior of the domain of an everywhere continuous function implies convexity on the whole domain is standard, but we show the main steps for completeness. Assume that $({\bm{x}}_{n})_{n\in{\mathbb{N}}}$ and $({\bm{y}}_{n})_{n\in{\mathbb{N}}}$ are two sequences in ${\mathbb{R}}_{++}^{N}$ converging in the normed vector space $({\mathbb{R}}^{N},\|\cdot\|)$ to, respectively, the vectors ${\bm{x}}\in{\mathbb{R}}_{+}^{N}$ and ${\bm{y}}\in{\mathbb{R}}_{+}^{N}$. Given $\alpha\in~]0,1[$, convexity of $g_{T_{\|\cdot\|}}$ on ${\mathbb{R}}_{++}^{N}$ implies that $(\forall n\in{\mathbb{N}})~g_{T_{\|\cdot\|}}(\alpha{\bm{x}}_{n}+(1-\alpha){\bm{y}}_{n})\leq\alpha g_{T_{\|\cdot\|}}({\bm{x}}_{n})+(1-\alpha)g_{T_{\|\cdot\|}}({\bm{y}}_{n})$. Passing to the limit as $n\to\infty$ and recalling that $g_{T_{\|\cdot\|}}$ is continuous on ${\mathbb{R}}^{N}_{+}$, we deduce $g_{T_{\|\cdot\|}}(\alpha{\bm{x}}+(1-\alpha){\bm{y}})\leq\alpha g_{T_{\|\cdot\|}}({\bm{x}})+(1-\alpha)g_{T_{\|\cdot\|}}({\bm{y}})$, which establishes convexity of $g_{T_{\|\cdot\|}}$ on ${\mathbb{R}}_{+}^{N}$. In particular, setting $\alpha$ to $\alpha=1/2$ and using homogeneity of $g_{T_{\|\cdot\|}}$, we deduce $\alpha g_{T_{\|\cdot\|}}({\bm{x}}+{\bm{y}})\leq\alpha g_{T_{\|\cdot\|}}({\bm{x}})+\alpha g_{T_{\|\cdot\|}}({\bm{y}})$, which proves that $g_{T_{\|\cdot\|}}$ satisfies the triangle inequality on ${\mathbb{R}}_{+}^{N}$. Therefore, $g_{T_{\|\cdot\|}}$ satisfies all properties of a norm when restricted to vectors in the closed cone ${\mathbb{R}}_{+}^{N}$, so the set $C_{T_{\|\cdot\|}}:=\{{\bm{x}}\in{\mathbb{R}}^{N}_{+}~\mid g_{T_{\|\cdot\|}}({\bm{x}})\leq 1\}$ is a compact convex set with nonempty interior. Furthermore, $C_{T_{\|\cdot\|}}$ is downward comprehensive on ${\mathbb{R}}_{+}^{N}$ as a consequence of Proposition 2(iii). All the properties of $C_{T_{\|\cdot\|}}$ we have proved enable us to apply [8, Proposition 2] to conclude that the Minkowski functional in (14) is indeed a monotone norm, and it satisfies
@@
The result in Proposition 3 provides a foundation for a natural definition of a norm as follows:
@@
Let $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ and ${T_{\|\cdot\|}}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}$ satisfy hypothesis ($T{T_{\|\cdot\|}}$) for a given monotone norm $\|\cdot\|$. If the function ${\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto\rho(\mathrm{diag}({\bm{x}}){T_{\|\cdot\|}})$ is convex on ${\mathbb{R}}_{++}^{N}$, we say that $T_{\|\cdot\|}$} is norm-inducing, and we call the norm $\|\cdot\|_{T_{\|\cdot\|}}$ in (14) the norm induced by $T_{\|\cdot\|}$}.
Our primary focus is on constructing norm-inducing general interference mappings derived from an explicitly given standard interference mapping and a monotone norm. However, in certain instances, it is straightforward to extract a suitable standard interference mapping and a monotone norm from a general interference mapping $G$ to verify that $G$ is norm-inducing, as illustrated in the next example, which is common in wireless networks – see, for example, the set in (28) in Sect. III-B.